3. For function y(t) with y(0) = 1, y(0) = -2, y(0) = 3, y(0) = 2020, find c{dey!
Answer #1

Detailed solutions is attached.please like itt-doo Paget. La since we know that Reig = 1 st Flat . 4play for inte plats at = Last Fa, 7,0 +8.1 28+ Fedt => {{ flesyz Lm es3= (ن الي ا= (90 2- = (ناو - عود ايمان او را هر کدام کیا وج = او - ار 3 - (ye 2 - اه وا}L في y 3- ( 2- (1) - ايه 29 L 33 = 2